Two dice each have faces numbered 1 to 6. They are rolled together. Choose the option that gives the probability that the difference of the numbers obtained is 1.
1
Expert's answer
2011-10-07T08:53:29-0400
Omega={(x1,x2)|x1=1,..,6; i=1,2}
A={(x1,x2)|x1-x2=1}U{(x1,x2)|x2-x1=1}
P(A)=|A| / |Omega| = (5+5)/(6*6)=10/36